6) 5.00m2ー? ft®

conversion in scientific form or decimal form. not both. please show work or explain so I will know how to do this.
Answer #1

1 ft = 0.3048 m

1m = 1/0.348 = 3.28 ft

1m^3 = (3.28)^3 ft^3

= 35.314 ft^3

5 m^3 = 5*35.314 = 176.573 ft^3